# Judging the Judges: A Systematic Study of Position Bias in LLM-as-a-Judge

### v7

@@ Abstract
LLM-as-a-Judge has emerged as a promising alternative to human evaluators across various tasks, yet inherent biases—particularly position bias, the tendency to favor solutions based on their position within the prompt—compromise its reliability. This study investigates position bias in LLM judges across pairwise and list-wise comparison settings, introducing three metrics: repetition stability, position consistency, and preference fairness. Our experiments, involving 12 LLM judges across MTBench and DevBench with 22 tasks and approximately 40 solution-generating models, result in over 100,000 evaluation instances. We identify Judge-Level, Candidate-Level, and Task-Level factors contributing to bias. The findings confirm that position bias is not due to random chance and varies significantly across judges and tasks. While position bias is weakly influenced by the length of prompt components, it is strongly affected by the quality gap between solutions. Our agreement and disagreement analysis among judges further provides insights into the distribution of judging difficulty across the dataset, and highlights the potential for dataset modifications.
@@
Lin Shi, Chiyu Ma, Wenhua Liang, Weicheng Ma, Soroush Vosoughi Dartmouth College {lin.shi.26, chiyu.ma.gr, wendy.liang.25, weicheng.ma.gr, soroush.vosoughi}@dartmouth.edu
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have emerged as evolutionary technologies, gathering global interest and stimulating substantial research into their applications. Evaluating LLMs has received increasing attention due to their advancing capabilities across diverse fields. While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability and reproducibility in extensive evaluations Zeng et al. (2023); Karpinska et al. (2021). To automate evaluations and reduce reliance on costly human evaluators, the LLM-as-a-Judge methodology emerged as a promising alternative across various tasks. Despite a high level of agreement with human judgments Zheng et al. (2024b); Li et al. (2024a); Zhu et al. (2023), inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
Position bias refers to the tendency of LLM judges to favor certain positions within prompt components rather than the content itself, as shown in Fig. 1 (a). This bias has been prevalent in all types of LLM judges Qin et al. (2024); Li et al. (2023d), raising concerns about their reliability. Previous studies Zheng et al. (2024a, b); Zeng et al. (2023) have identified position bias alongside other biases and assessed its impact. Although mitigation strategies have been proposed, they often fall short due to incomplete bias elimination Guo et al. (2024), increased complexity Li et al. (2024b); Khan et al. (2024); Chua et al. (2024), the introduction of new biases Ohi et al. (2024), or inconsistent effectiveness Gallegos et al. (2024). These persistent challenges highlight the limitations of current approaches and point to a possible incomplete understanding of position bias. Although numerous studies have evaluated position bias in LLM-as-a-Judge settings, the underlying factors contributing to this phenomenon remain insufficiently explored.
In this study, we provide an in-depth and systematic investigation into position bias within the context of LLM-as-a-Judge. While evaluating Position Consistency, one of the most widely used methods for measuring position bias, we introduce two novel metrics: Preference Fairness and Repetition Stability. Specifically, we move beyond simply assessing Position Consistency by incorporating Preference Fairness, which provides deeper insights into the specific answer directions where models exhibit unfair preferences. Additionally, the measurement of Repetition Stability ensures that the observed position bias in the given model and tasks is not due to random variations, thus strengthening the reliability of the findings.
@@ 2 Evaluation Settings & Definitions
We begin by outlining the settings for pairwise and list-wise comparisons employed in our experiments for LLM-as-a-Judge. Following this, we define the three metrics used in our evaluation: Position Consistency (PC), Preference Fairness (PF), and Repetition Stability (RS). Finally, we provide a detailed description of the factors we found that are related to position bias at the Judge-Level, Candidate-Level, and Task-Level.
@@ 2.1 Pairwise & List-wise Comparison
Pairwise Comparison: In the context of pairwise comparison, LLM judges are tasked with selecting the better solution provided by two candidate models in response to a given task question. As shown in Fig. 1 (a), the system prompt, option choices, task question, and solutions from two candidate models (original prompt) are presented to the LLM judges to select the better solution. The experiment is conducted in a double-blind setting. The identities of the candidate models are hidden from the LLM judges, and the candidate models are unaware that their solutions will be compared to another model when answering the question. Then, the prompt with solutions in a swapped position (swapped prompt) is given to the same judge again, which results in a judgment pair. If the LLM judge consistently favors the same solution regardless of the swapped position, it is considered position consistent. Conversely, if the LLM judge selects different winners, position bias is observed, with the preference direction being either primacy (e.g. always choose $\{A\}$) or recency (e.g. always choose $\{B\}$). Example of measuring preference fairness with specific choice pairs is shown in Fig. 1 (b). To accommodate the possibility of ties, various option modes are employed: Two-Option mode restricts judges to choosing between two options, labeled A for the first candidate and B for the second. Three-Option mode introduces an additional choice, C, allowing judges to indicate a tie if neither solution is preferable, as illustrated in Fig. 1 (a). These option modes were explicitly specified in the system prompts to ensure clear guidance for the decision-making process of the LLM judges.
List-wise Comparison: Unlike pairwise settings, where LLM judges select the superior solution from two candidates, list-wise comparative approaches involve evaluating three or more candidates simultaneously, as shown in Fig. 1 (c). For efficiency, we prompt judges to select the best candidate rather than ranking the entire list. The “swapped setting" used in pairwise evaluations is generalized to order permutations for list-wise judgments, ensuring that each candidate appears in every possible position exactly once. For a list of $p$ candidates, this results in $p$ permutations. In the $i^{th}$ permutation, the $i^{th}$ candidate is set to appear in the first position. Additionally, an option is provided to account for ties, allowing judges to indicate if there is no certainly superior solution.
@@ 2.2 Evaluation Metrics
In our study, we first verify if the capable LLMs with a high repetition stability and then evaluate their position bias in terms of position consistency and preference fairness. The metrics are introduced as follows:
@@
Position Consistency ($PC$) quantifies how frequently a judge model prefers the same solution after the order of solutions is permuted. It is calculated as the ratio of consistent evaluation series to the total number of valid evaluations, where a series is deemed consistent if the judge model prefers the same winning solution across permutations. Formally, it is calculated as
@@
where $V$ is the set of choices that correspond to position consistency, and $(C_{1}^{j},\dots,C_{p}^{j})$ denotes the judgment series for the $j^{th}$ query when there are $p$ candidate solutions in the list, and $n$ represents the number of prompt series. An example of such series of choices under pairwise comparison setting can be found in Fig. 1 (b). This formula aims to provide a direct measure of a judge model’s position bias and has been widely used in previous studies for its simplicity.
Preference Fairness ($PF$) measures the extent to which judge models favor certain solution positions. In pairwise comparisons, an LLM judge may exhibit a preference for either primacy or recency. These terms replace the more verbose “preference for the first/second candidate model" used in previous studies, ensuring clarity and generalization for future research. The examples of such preferences are demonstrated in Fig. 1 (b). Previous studies proposed two common ways to measure the preference fairness. One way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number ($pcn$) and recency-count-number ($rcn$). The counts are then normalized by the total number of prompt pairs (Zheng et al., 2024b; Zhu et al., 2023). However, the sensitivity of this measurement highly depends on the size of dataset, making comparisons across datasets unreliable, especially when the number of questions and instances varies for each task.
@@ 2.3 Factors Affecting Position Bias
Table 1: Factors influencing position bias. Significant factors, identified via bidirectional stepwise regression, are marked with * and highlighted in red based on empirical findings. Task Input refers to the question itself, while Task Output denotes the candidate model’s answers, serving as both Candidate-level and Task-level factors. Prompt includes the full query presented to judges: Task Input, Task Output, and system prompts.
Among the influencing factors, we selected “familial property” for Judge-level factors, as it reflects similar model sizes and training specifics, which are often proprietary and not publicly accessible. The familial categories of the models used in our studies are (1) GPT, (2) Claude, and (3) Gemini allowing for straightforward grouping by company and version. More details and discussions about the familial property can be found in Appendix Sec. C.
Answer quality gap, based on our findings, is highly influential yet under-explored in prior research. We define the quality of a candidate’s solution by how effectively it addresses the question. Consequently, the answer quality gap refers to the disparity in quality between the solutions from one candidate model and the others to the same question and hence considered the Candidate-level factor. Ideally, when a reliable LLM judge is presented with a question and corresponding answer pairs or series, it would prefer the highest-quality answer, where the corresponding candidate is denoted as the winner selected by the judge.
3 Experiment
3.1 Experiment Settings
In this study, we evaluated twelve up-to-date close-sourced models from the GPT (OpenAI, 2023), Claude (Anthropic, 2024), and Gemini (Gemini Team, 2024) series using our framework. Additionally, three Llama models (Touvron et al., 2023), with parameters ranging from 7B to 13B, were selected as open-source exemplars due to their popularity and recognized capability in reasoning tasks. However, due to their limited context window size and failure to follow the output format, which led to suboptimal performance on our tasks (as shown in Table 2), we focused our investigation exclusively on more powerful closed-source models. We included the results of open-source models for completeness. More discussions can be found in Appendix.Sec. D
We adopted the modified MTBench (Zheng et al., 2024b) and DevBench (Li et al., 2024a) datasets for our study due to their demonstrated high human-LLM agreement and the reliability of state-of-the-art LLMs on the evaluation tasks. For pairwise comparisons, We fixed one of the candidates as vicuna-13b-v1.3 for MTBench and human for DevBench to serve as baselines, ensuring decent quality of responses. MTBench consists of 30 candidate models, 8 tasks, and 10 questions per task; for DevBench, we divide the general metric into more detailed ones and consider them as different tasks, resulting in 10 candidate models, 14 tasks, and 8 questions per task. We then paired solutions of these candidate models with that of the baseline candidate for evaluation by the LLM judges. We adopted Two-option mode for MTBench, and Three-option mode for DevBench. For list-wise experiments, we randomly sampled 9 models to form three triple-candidate lists and evaluated four representative judges on MTBench. The prompt templates we used are identical to those in the benchmarks for pairwise comparisons, with minor modifications to accommodate list-wise evaluations. More details about the models, tasks, and prompts can be found in Appendix. Sec. G.
To compute repetition stability, we sampled 3 questions per task and 4 candidate models, paired with baseline candidates, for each judge to evaluate across 3 repetitive trials. This resulted in 576 instances per judge for MTBench and 432 instances per judge for DevBench. The temperature hyperparameter was set to 1 for all judge models to generate nontrivial results. To compute position consistency and preference fairness, the number of instances increased to 4,800 and 2,240, covering the entire MTBench and DevBench datasets. In total, more than 100,000 evaluation instances were analyzed in this study.
To identify significant factors contributing to position bias, we performed bidirectional stepwise regression on data from the two benchmarks. We used variables such as average lengths of input, output, and prompt; answer quality gap; LLM judge series; candidate identities; and task categories to predict $PC$ and $PF$, respectively. Each model prunes non-significant variables based on the Akaike Information Criterion (AIC) score. This process involves both forward selection and backward elimination, with each "step" testing whether including or excluding a variable improves the model’s AIC value. Further details about the process can be found in Appendix. Sec. F.
@@ 3.2 Empirical Results
The evaluation results of 12 close-source and 3 open-source models in terms of repetition stability, position consistency, and preference fairness on MTBench and DevBench are listed in Table 2. For each judge, we calculate its average $RS$, $PC$, and $PF$ across all candidates and tasks. For $RS$ and $PC$, higher values are preferable. A high $RS$ value is particularly important as a prerequisite for meaningful computations of $PC$ and $PF$, ensuring the LLM judge’s choice patterns are not due to random variations. Fig. 2 (a)(b) demonstrate that position bias varies by judges and tasks significantly. Fig. 2 (c) explores the correlation between the metrics $PC$ and $PF$. Fig. (d) to (f) further investigate the impact of the answer quality gap on position bias. These analyses were conducted by considering all judges together on MTBench. The detailed analyses for each judge and on DevBench can be found in Appendix.Sec.E.
Through bidirectional stepwise regression, we found that, based on data from the two benchmarks, LLM judge series, candidate identities, and task categories significantly impact Position Consistency among all variables. Similarly, these factors also contribute significantly to Preference Fairness. Additionally, we found that average output length is a statistically significant predictor of $PF$. This finding is not surprising, as longer outputs are generally perceived as higher quality and more preferred. Quantitative results and more discussions can be found in Appendix. Sec. F.
@@
Llama-2-7B 0.83 $\pm$ 0.15 0.44 $\pm$ 0.19 0.11 0.17 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00
Llama-2-13B 0.45 $\pm$ 0.22 0.56 $\pm$ 0.33 -0.08 0.55 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00
Llama-3-8B 0.16 $\pm$ 0.18 0.43 $\pm$ 0.39 -0.05 0.84 0.06 $\pm$ 0.09 0.04 $\pm$ 0.09 -0.05 0.94
Table 2: Evaluation results for Repetition Stability ($RS$), Position Consistency ($PC$), and Preference Fairness ($PF$) are presented for both pairwise and list-wise evaluation approaches, with the top 5 performances marked in bold. Errors arise from judgment failures (e.g., exceeding context window, not following output format). High error rates and low $RS$ are marked red, rendering further evaluations invalid due to insufficient data. Open-source models (e.g., Llama) are excluded due to high error rates with our lengthy instances (See Appendix.Sec.D for more discussions). List-wise evaluation is conducted on four representative judges from each family to validate scalability.
Figure 2: Judge performances on MTBench. Fig. (a)(b) are the radar charts for the $PC$ comparison by family, judge, and task Fig. (c) to (f) investigate the impact of task output length and answer quality gap on position bias. Fig. (a)(c)(d)(e) are for pairwise comparative settings, while Fig. (b)(f) are obtained under list-wise evaluations.

### v8

@@ Abstract
LLM-as-a-Judge has emerged as a promising alternative to human evaluators across various tasks, yet inherent biases—particularly position bias, the tendency to favor solutions based on their position within the prompt—compromise its reliability. This exploratory study evaluates position bias in LLM judges across pairwise and list-wise comparison settings, introducing three metrics: repetition stability, position consistency, and preference fairness. Our experiments, involving 15 LLM judges across MTBench and DevBench with 22 tasks and approximately 40 solution-generating models, result in over 150,000 evaluation instances. We identify Judge-Level, Candidate-Level, and Task-Level factors contributing to bias. The findings confirm that position bias is not due to random chance and varies significantly across judges and tasks. While position bias is weakly influenced by the length of prompt components, it is strongly affected by the quality gap between solutions. Our agreement and disagreement analysis among judges further provides insights into the distribution of judging difficulty across the dataset, and highlights the potential for dataset modifications.
@@
Lin Shi, Chiyu Ma, Wenhua Liang, Xingjian Diao, Weicheng Ma, Soroush Vosoughi Dartmouth College {lin.shi.26, chiyu.ma.gr, wendy.liang.25, xingjian.diao.gr, weicheng.ma.gr, soroush.vosoughi}@dartmouth.edu
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have emerged as evolutionary technologies, gathering global interest and stimulating substantial research into their applications. Evaluating LLMs has received increasing attention due to their advancing capabilities across diverse fields. While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability in extensive evaluations Zeng et al. (2023); Karpinska et al. (2021). To automate evaluations and reduce reliance on costly human evaluators, the LLM-as-a-Judge methodology emerged as a promising alternative across various tasks. Despite a high level of agreement with human judgments Zheng et al. (2024b); Li et al. (2024a); Zhu et al. (2023), inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
Position bias refers to the tendency of LLM judges to favor certain positions within prompt components rather than the content itself, as shown in Fig. 1 (a). This bias has been prevalent in all types of LLM judges Qin et al. (2024); Li et al. (2023d), raising concerns about their reliability. Previous studies Zheng et al. (2024a, b); Zeng et al. (2023) have identified position bias alongside other biases and assessed its impact. Although mitigation strategies have been proposed, they often fall short due to incomplete bias elimination Guo et al. (2024), increased complexity Li et al. (2024b); Khan et al. (2024); Chua et al. (2024), the introduction of new biases Ohi et al. (2024), inconsistent effectiveness Gallegos et al. (2024), or impracticality for closed-source models, as mechanistic solutions require modifications to model internals Wang et al. (2025); Yu et al. (2025). From an exploratory perspectives, despite empirical studies evaluating position bias in pairwise LLM-as-a-Judge settings Zheng et al. (2024b); Wang et al. (2023), the underlying factors contributing to this phenomenon and extended measurements in list-wise comparative paradigms remain incompletely explored.
In this study, we provide an in-depth and systematic investigation into position bias within the context of LLM-as-a-Judge. While evaluating Position Consistency Zheng et al. (2024b), one of the most widely used methods for measuring position bias, we introduce two novel metrics: Preference Fairness and Repetition Stability. Specifically, we move beyond simply assessing Position Consistency by incorporating Preference Fairness, which provides deeper insights into the specific answer directions where models exhibit unfair preferences. Additionally, the measurement of Repetition Stability ensures that the observed position bias in the given model and tasks is not due to random variations, thus strengthening the reliability of the findings.
@@ 2 Evaluation Settings & Definitions
We begin by outlining the settings for pairwise and list-wise comparisons employed in our experiments for LLM-as-a-Judge. Following this, we define the three metrics used in our evaluation: Position Consistency (PC), Preference Fairness (PF), and Repetition Stability (RS). Finally, we provide a detailed description of the factors we found that are related to position bias at the Judge-Level, Candidate-Level, and Task-Level. Our exploratory study was conducted post hoc, meaning the LLM judgments were collected first, and the factors influencing position bias were then identified and analyzed.
@@ 2.1 Pairwise & List-wise Comparison
Pairwise Comparison: In the context of pairwise comparison, LLM judges are tasked with selecting the better solution provided by two candidate models in response to a given task question. As shown in Fig. 1 (a), the system prompt, option choices, task question, and solutions from two candidate models (original prompt) are presented to the LLM judges to select the better solution. The experiment is conducted in a double-blind setting. The identities of the candidate models are hidden from the LLM judges, and the candidate models are unaware that their solutions will be compared to another model when answering the question. Then, the prompt with solutions in a swapped position (swapped prompt) is given to the same LLM judge again, which results in a judgment pair. If the LLM judge consistently favors the same solution regardless of the swapped position, it is considered position consistent. Conversely, if the LLM judge selects different winners, position bias is observed, with the preference direction being either primacy (e.g. always choose $\{A\}$) or recency (e.g. always choose $\{B\}$). Example of measuring preference fairness with specific choice pairs is shown in Fig. 1 (b). To accommodate the possibility of ties, various option modes are employed: Two-Option mode restricts LLM judges to choosing between two options, labeled A for the first candidate and B for the second. Three-Option mode introduces an additional choice, C, allowing LLM judges to indicate a tie if neither solution is preferable, as illustrated in Fig. 1 (a). These option modes were explicitly specified in the system prompts to ensure clear guidance for the decision-making process of the LLM judges.
List-wise Comparison: Unlike pairwise settings, where LLM judges select the superior solution from two candidates, list-wise comparative approaches involve evaluating three or more candidates simultaneously, as shown in Fig. 1 (c). For efficiency, we prompt LLM judges to select the best candidate rather than ranking the entire list. The “swapped setting" used in pairwise evaluations is generalized to order permutations for list-wise judgments, ensuring that each candidate appears in every possible position exactly once. For a list of $p$ candidates, this results in $p$ permutations. In the $i^{th}$ permutation, the $i^{th}$ candidate is set to appear in the first position. Additionally, an option is provided to account for ties, allowing LLM judges to indicate if there is no certainly superior solution.
@@ 2.2 Evaluation Metrics
In our study, we first verify whether capable LLM judges exhibit high repetition stability and then evaluate their position bias in terms of position consistency and preference fairness. The metrics are introduced as follows:
@@
Position Consistency ($PC$) quantifies how frequently LLM judges prefer the same solution after the order of solutions is permuted. It is calculated as the ratio of consistent evaluation series to the total number of valid evaluations, where a series is deemed consistent if the LLM judge prefers the same winning solution across permutations. Formally, it is calculated as
@@
where $V$ is the set of choices that correspond to position consistency, and $(C_{1}^{j},\dots,C_{p}^{j})$ denotes the judgment series for the $j^{th}$ query when there are $p$ candidate solutions in the list, and $n$ represents the number of prompt series. An example of such series of choices under pairwise comparison setting can be found in Fig. 1 (b). This formula aims to provide a direct measure of a LLM judge’s position bias and has been widely used in previous studies for its simplicity.
Preference Fairness ($PF$) measures the extent to which LLM judges favor certain solution positions. In pairwise comparisons, an LLM judge may exhibit a preference for either primacy or recency. These terms replace the more verbose “preference/bias for the first/second candidate model" used in previous studies Zheng et al. (2024b), ensuring clarity and generalization for future research. The examples of such preferences are demonstrated in Fig. 1 (b). Previous studies proposed two common ways to measure the preference fairness. One way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number ($pcn$) and recency-count-number ($rcn$). The counts are then normalized by the total number of prompt pairs (Zheng et al., 2024b; Zhu et al., 2023). However, the sensitivity of this measurement highly depends on the size of dataset, making comparisons across datasets unreliable, especially when the number of questions and instances varies for each task.
@@ 2.3 Factors Affecting Position Bias
Among the influencing factors, we selected “familial property” for Judge-level factors, as it reflects similar model sizes or training specifics, which are often proprietary and not publicly accessible for closed-source capable models. The familial categories of the models used in our studies are (1) GPT, (2) Claude, (3) Gemini, and (4) Llama allowing for straightforward grouping by company and version. More details and discussions about the familial property can be found in Appendix Sec. C.
Answer quality gap: While prior studies Wang et al. (2023) explored quality disparities using "score gaps" in score-based LLM-as-a-Judge, this factor remains under-explored in comparative settings, which we address by introducing "answer quality gap" for both pairwise and list-wise evaluations. We define the quality of a candidate’s solution by how effectively it addresses the question. Consequently, the answer quality gap refers to the disparity in quality between the solutions from one candidate model and the others to the same question and hence considered the Candidate-level factor. Ideally, when a reliable LLM judge is presented with a question and corresponding answer pairs or series, it would prefer the highest-quality answer, where the corresponding candidate is denoted as the winner selected by the LLM judge.
3 Experiment
3.1 Experiment Settings
In this study, we evaluated position bias of 15 models from the GPT (OpenAI, 2023), Claude (Anthropic, 2024), Gemini (Gemini Team, 2024), and Llama (Touvron et al., 2023) series using our framework. For datasets, we adopted the modified MTBench (Zheng et al., 2024b) and DevBench (Li et al., 2024a) due to their demonstrated high human-LLM agreement and the validated reliability of state-of-the-art LLM judges on the evaluation tasks. For pairwise comparisons, We fixed one of the candidates as vicuna-13b-v1.3 for MTBench and human for DevBench to serve as baselines, ensuring decent quality of solutions to the given questions. MTBench consists of 30 candidate models, 8 tasks, and 10 questions per task; for DevBench, we divide the general metric into more detailed ones and consider them as different tasks, resulting in 10 candidate models, 14 tasks, and 8 questions per task. We then paired solutions of these candidate models with that of the baseline candidate for evaluation by the LLM judges.
We adopted Two-option mode for MTBench, and Three-option mode for DevBench. For list-wise experiments, we randomly sampled 9 models to form three triple-candidate lists and evaluated four representative LLM judges on MTBench. The prompt templates we used are identical to those in the benchmarks for pairwise comparisons, with minor modifications to accommodate list-wise evaluations. More details about the models, tasks, and prompts can be found in Appendix. Sec. F.
To compute repetition stability, we sampled 3 questions per task and 4 candidate models, paired with baseline candidates, for each LLM judge to evaluate across 3 repetitive trials. This resulted in 576 instances per judge for MTBench and 432 instances per judge for DevBench. The temperature hyperparameter was set to 1 for all LLM judges to generate nontrivial results. To compute position consistency and preference fairness, the number of instances increased to 4,800 and 2,240, covering the entire MTBench and DevBench datasets. In total, more than 100,000 evaluation instances were analyzed in this study.
To identify significant factors contributing to position bias, we performed bidirectional stepwise regression on data from the two benchmarks. We used variables such as average lengths of input, output, and prompt; answer quality gap; LLM judge series; candidate identities; and task categories to predict $PC$ and $PF$, respectively. Each model prunes non-significant variables based on the Akaike Information Criterion (AIC) score. This process involves both forward selection and backward elimination, with each "step" testing whether including or excluding a variable improves the model’s AIC value. Further details about the process can be found in Appendix. Sec. E.
@@ 3.1 Experiment Settings
Table 1: Factors influencing position bias. Significant factors, identified via bidirectional stepwise regression, are marked with * and highlighted in red based on empirical findings on both MTBench and DevBench results. Task Input refers to the question itself, while Task Output denotes the candidate model’s answers, serving as both Candidate-level and Task-level factors. Prompt includes the full query presented to LLM judges: Task Input, Task Output, and system prompts.
@@ 3.2 Empirical Results
The evaluation results of 12 close-source and 3 open-source LLM judges in terms of repetition stability, position consistency, and preference fairness on MTBench and DevBench are listed in Table 2. For each judge, we calculate its average $RS$, $PC$, and $PF$ across all candidates and tasks. For $RS$ and $PC$, higher values are preferable. A high $RS$ value is particularly important as a prerequisite for meaningful computations of $PC$ and $PF$, ensuring the LLM judge’s choice patterns are not due to random variations. Fig. 2 (a)(b) demonstrate that position bias varies by judges and tasks significantly. Fig. 2 (c) explores the correlation between the metrics $PC$ and $PF$. Fig. (d) to (f) further investigate the impact of the answer quality gap on position bias. These analyses were conducted by considering all judges together on MTBench. More analyses can be found in Appendix. Sec. D.
Through bidirectional stepwise regression, as shown in Table 1, LLM judge series, candidate identities, and task categories significantly impact Position Consistency among all variables. Similarly, these factors also contribute significantly to Preference Fairness. Additionally, we found that average output length is a statistically significant predictor of $PF$. This finding is not surprising, as longer outputs are generally perceived as higher quality and more preferred. Quantitative results and more discussions can be found in Appendix. Sec. E.
@@
Llama-3.3-70B 0.96 $\pm$ 0.06 0.80 $\pm$ 0.16 -0.05 0.00 0.99 $\pm$ 0.01 0.89 $\pm$ 0.12 -0.03 0.00
Llama-3.1-405B 0.93 $\pm$ 0.10 0.77 $\pm$ 0.16 0.10 0.02 0.94 $\pm$ 0.10 0.79 $\pm$ 0.15 0.01 0.00
Llama-3.1-8B 0.75 $\pm$ 0.32 0.69 $\pm$ 0.23 -0.03 0.25 0.79 $\pm$ 0.36 0.47 $\pm$ 0.18 0.25 0.00
Table 2: Evaluation results for Repetition Stability ($RS$), Position Consistency ($PC$), and Preference Fairness ($PF$) are presented for both pairwise and list-wise evaluation approaches, with the top 5 performances marked in bold. Errors arise from judgment failures (e.g., exceeding context window, not following output format). High error rates and low $RS$ are marked red, rendering further evaluations invalid due to insufficient data. List-wise evaluation is conducted on four representative judges to validate scalability.